OpenStudy (anonymous):

medal!! and become fan!! One day, a person went to a horse racing area. Instead of counting the number of humans an horses, he conuted 74 heads and 196 legs. How many humans and horses were there? A. 37 humans and 98 horses B. 24 horses and 50 humans C. 31 horses and 74 humans D.24 humans and 50 horses

OpenStudy (anonymous):

I think heads are for horses too

OpenStudy (amoodarya):

head of human =x leg of human =2x horse head=y horse leg=4y so \[x+y=74\\2x+4y=196\] now solve the system of equation

OpenStudy (anonymous):

no ... if they are 196 legs thts not 74 humans

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok.. i did the equation and i end up with B

OpenStudy (amoodarya):

if you solve the system y must b 24 and x=74-24 so horses are 24 and ...

OpenStudy (anonymous):

human are 50!!!

OpenStudy (amoodarya):

yes